4.6m + 9.2 = 3.4m + 19.4

Answer by avani(25)

4.6m + 9.2 = 3.4m + 19.4
4.6m-3.4m=19.4-9.2
1.2m=10.20
m=10.20/1.2=8.50
Thus, m=8.50